A current value of $50,000 is held in an account earning 12% APR. Equal payments are to be paid out quarterly until the account is depleted after 8 years. How much would the equal quarterly payments be under each of the following interest compounding scenarios? (Hint: first find the effective interest for each scenario to the nearest 10,000th of a percent. Then, find the value of the equal payments to the nearest $0.01.)

A. 12% APR Compounded Quarterly (4 quarters per year)

B. 12% APR Compounded Monthly (3 months per quarter)

C. 12% APR Compounded Daily (90 days per quarter)

D. 12% APR Compounded Continuously

please answer will rate
Answer #1

1.
=50000*(12%/4)/(1-1/(1+12%/4)^(4*8))=2452.33091377954

2.
=50000*(((1+12%/12)^(12/4)-1))/(1-1/(1+((1+12%/12)^(12/4)-1))^(4*8))=2462.38674020653

3.
=50000*(((1+12%/360)^(360/4)-1))/(1-1/(1+((1+12%/360)^(360/4)-1))^(4*8))=2467.35168473336

4.
=50000*((EXP(12%/4)-1))/(1-1/(1+(EXP(12%/4)-1))^(4*8))=2467.52413490153